Recent Log Entries

iocsl25rfhtr restored

Log Number: 3578612Logbook(s): Tags: Body: restorelib 1.0 Target restore date was 2018-06-22 14:44:18 The output from the myRestore execution was as follows: iocsl25rfhtr <NOTE> Rf12Htr_SL25x: Provided macro *IOC* is never used Pause: 15 seconds Rf12Htr_SL25x Rf12Htr_Save_SL25 56(56) restored Pause: 5 seconds Execute: caput R2YXHTSWI 1 (success) Execute: caput R2YXHTPD 55 (success) Execute: caput R2PXHTMAKE 1 (success) system 3(3) restored Execute: /usr/csite/certified/bin/ced2Epics -ced ced -A -s -i iocsl25rfhtr (success)

iocsl25rfhtr restored

Log Number: 3578607Logbook(s): Tags: Body: restorelib 1.0 Target restore date was 2018-06-22 14:28:01 The output from the myRestore execution was as follows: iocsl25rfhtr <NOTE> Rf12Htr_SL25x: Provided macro *IOC* is never used Pause: 15 seconds Rf12Htr_SL25x Rf12Htr_Save_SL25 56(56) restored Pause: 5 seconds Execute: caput R2YXHTSWI 1 (success) Execute: caput R2YXHTPD 55 (success) Execute: caput R2PXHTMAKE 1 (success) system 3(3) restored Execute: /usr/csite/certified/bin/ced2Epics -ced ced -A -s -i iocsl25rfhtr (success)

COMMENT ON ^Remove a Radiabeam Corrector Magnet from the GTS over to the Main Injector to be Installed

Log Number: 3578606Logbook(s): Body: 

We have complete the installation of the Radiabeam Corrector
Magnet in the Injector.

We installed (2) 10 ohm 50w resistors in parallel with the
(2) 15 ohm 25w resistors that were already in the junction
box, to give us a 6 ohm limited resistance on the magnets.

We cycle both MCR2I01H and MCR2I01V magnet and did a
polarity check.
Scott Higgins has updated the software for these magnet in
the CED and EPIC.

C.Humphry, L.Broeker, E.Diggs

COMMENT ON Test Operate The Spare IOT Tube

Log Number: 3578602Logbook(s): Body: 

Larry and I took the tube from #1 IOT (Pass 2) and replaced
it with the spare. Today we attempted to operate the IOT
but discovered problems with one of IOT 4's two bypass
switchs and problems with loose and damaged wires in the
bypass relay logic. One bad switch is replaced and the
loose wires are fixed. We plan to attempt to operate/test
IOT 1 with the spare tube again on Monday. This test will
not be induce RF into a cavity but into a load.

iocsl25rfhtr restored

Log Number: 3578601Logbook(s): Tags: Body: restorelib 1.0 Target restore date was 2018-06-22 13:40:53 The output from the myRestore execution was as follows: iocsl25rfhtr <NOTE> Rf12Htr_SL25x: Provided macro *IOC* is never used Pause: 15 seconds Rf12Htr_SL25x Rf12Htr_Save_SL25 56(56) restored Pause: 5 seconds Execute: caput R2YXHTSWI 1 (success) Execute: caput R2YXHTPD 55 (success) Execute: caput R2PXHTMAKE 1 (success) system 3(3) restored Execute: /usr/csite/certified/bin/ced2Epics -ced ced -A -s -i iocsl25rfhtr (success)

iocsl25rfhtr restored

Log Number: 3578596Logbook(s): Tags: Body: restorelib 1.0 Target restore date was 2018-06-22 13:20:47 The output from the myRestore execution was as follows: iocsl25rfhtr <NOTE> Rf12Htr_SL25x: Provided macro *IOC* is never used Pause: 15 seconds Rf12Htr_SL25x Rf12Htr_Save_SL25 56(56) restored Pause: 5 seconds Execute: caput R2YXHTSWI 1 (success) Execute: caput R2YXHTPD 55 (success) Execute: caput R2PXHTMAKE 1 (success) system 3(3) restored Execute: /usr/csite/certified/bin/ced2Epics -ced ced -A -s -i iocsl25rfhtr (success)

Gun Deca Rad signals duringHV conditioning

Log Number: 3578595Logbook(s): Entry Makers: freybergBody: 

There are 10 decarad monitors around the new Gun. Not sure of the calibration, edm screen says "Counts/sec".

highest levels are observed on IDRINJRAD02, approaching 500 cnts/s

iocin1mag restored

Log Number: 3578594Logbook(s): Tags: Body: restorelib 1.0 Target restore date was 2018-06-22 13:07:58 The output from the myRestore execution was as follows: iocin1mag magnet_MAG 228(228) restored Pause: 2 seconds magnet_HVAL 113(113) restored Pause: 2 seconds magnet_LOOP 113(113) restored Pause: 2 seconds magnet_CMPQ 226(226) restored in1mag 11(11) restored system 3(3) restored Execute: /usr/csite/certified/bin/ced2Epics -ced ced -A -s -i iocin1mag (success)

LCW system update

Log Number: 3578593Logbook(s): Entry Makers: jonescBody: 

All Accelerator LCW plants are now on Automatic and regulating. The pump code has been updated so that if a P/DP sensor fails, the pump will go to 98% output (as opposed to ramping down).

Function Generator for 30hz switcher system found in bad state

Log Number: 3578592Logbook(s): Body: 

Just doing a daily check on the Function Generator for the 30hz Switcher System and I found it in a bad state. It looked as if it had been power cycled (viewed from the Expert Screen). I took a ride over to the MCC and took a look at the unit. It was in remote mode but something just did not look right. I turned the GPIB controller off line and then power cycled the Function Generator and set it up manually with 60hz Freg, 0vdc output and Square wave selected for output waveform. I then turned the GPIb controller back on line and waited for the PC104 device to boot up and make connection.

iocsl25rfhtr restored

Log Number: 3578591Logbook(s): Tags: Body: restorelib 1.0 Target restore date was 2018-06-22 12:45:12 The output from the myRestore execution was as follows: iocsl25rfhtr <NOTE> Rf12Htr_SL25x: Provided macro *IOC* is never used Pause: 15 seconds Rf12Htr_SL25x Rf12Htr_Save_SL25 56(56) restored Pause: 5 seconds Execute: caput R2YXHTSWI 1 (success) Execute: caput R2YXHTPD 55 (success) Execute: caput R2PXHTMAKE 1 (success) system 3(3) restored Execute: /usr/csite/certified/bin/ced2Epics -ced ced -A -s -i iocsl25rfhtr (success)

Installation of Jumpers to allow RF permits for replacement Cryomodule 1L07

Log Number: 3578589Logbook(s): Body: 

Sam, Leo and I installed the necessary jumpers in the Cross Connects @ 1L07B12 and 1L07B14 equipment raacks to allow RF to be generated in the replacement Cryomodule.
We will continue check out Monday..
It is possible that the diagrams we are using has a typo on one of the locations we will verify tghis Monday.
Thanks Leo and Sam for the help!

COMMENT ON Vacuum Activities SAD 2018

Log Number: 3578587Logbook(s): Body: 

Girder's 1L23 and 1L24 Leak check = Good with No leaks.
VIP's are turned On will pump with Turbo until ~1 PM, then
we will valve out Turbo from Girder then start the breakdown
of our equipment for movement to Cryomodule 1L07.

Follow-up Re: CHL2 Cooling Tower vulnerable status

Log Number: 3578585Logbook(s): Body: 

CHL Cooling Water Pump Motor 3 requires a replacement of the power cable. Facilities is preparing to repair or replace Pump 4 in addition to replacing cable for Pump Motor 3.
The differences in pump size, motor size, and electrical power requirements. A CHL1 pump can’t be placed into CHL2 location.

Pages